a gemologist determines the index of refraction of a crystal. if the angle of incidence in air is 50 degrees, the angle of refraction in the crystal is 20 degrees what is the index of refraction of the crystal?


1
Expert's answer
2021-06-03T09:17:36-0400

"n=\frac{\sin\alpha}{\sin\beta}=\frac{\sin50°}{\sin 20°}=2.24."